A rectangular field measures 56 m by 89 m. How much shorter is it to walk diagonally across the field than along two adjoining sides?
Answer by fractalier(6550)

Find the hypotenuse via
a^2 + b^2 = c^2
and subtract that from 56+89...thus we get
c^2 = 56^2 + 89^2 and
c = 105.15 m
145 - 105.15 = 39.85 meters saved